OpenStudy (saldis):

Graph ​ y+6=45(x+3)y+6=45(x+3) ​.

OpenStudy (518nad):

Graph ​ y+6=45(x+3),

OpenStudy (saldis):

Thanks babe...

OpenStudy (518nad):

do you know how to graph when its y=mx+b

OpenStudy (518nad):

xD when did i become ur babe

OpenStudy (saldis):

Idk since like idk two days ago xD

OpenStudy (518nad):

hello, so just find any 2 points on this line and join them together

OpenStudy (518nad):

what is y when x=0?

OpenStudy (saldis):

Ok

OpenStudy (518nad):

y+6=45(x+3), so plug x=0 in y+6=45*(0+3) y+6=45*3 y=45*3/6 =22.5 (0,22.5) is a point and -3,-6 is another point as this point makes the equation true too, check x=-3,y=-6

OpenStudy (518nad):



OpenStudy (518nad):